Good Morning: I haven't been able to solve this problem... I have to use the discriminant to determine the number of solutions of the quadratic equation, and whether the solutions are real or complex. Note: It is not necessary to find the roots; just determine the number and types of solutions.
2x^2 - 6x + 5 = 0 Answer by Prithwis(166) 
 Discriminant = (b^2 - 4ac) for ax^2+bx+c
Here a = 2, b=-6 and c = 5
b^2 - 4ac = 36-40 = -4
If b^2 - 4ac =0, there is only one real solution.
If b^2 - 4ac > 0, both the roots are real.
If b^2 - 4ac < 0, both the roots are non-real complex.
Here the discriminant is -4 (<0); so there are two non-real complex solutions of the given equation.